OpenStudy (anonymous):

Working together, Paul and Daniel can pick forty bushels of apples in 4.95 hours. Had he done it alone it would have taken Daniel 9 hours. Find how long it would take Paul to do it alone.

OpenStudy (cj49):

Paul and Daniel =1/4.95 of the job per hr

OpenStudy (cj49):

paul takes x hrs alone so he does 1/x of the job per hr

OpenStudy (cj49):

1/9 + 1/x = 1/4.95

OpenStudy (anonymous):

Thanks bae <3

OpenStudy (cj49):

x+9/9x=1/4.95

OpenStudy (cj49):

-4.05x=-44.55 x=44.55/4.05 x=11

OpenStudy (cj49):

^.^ np

OpenStudy (anonymous):

daniel collected 40 bushels of apples in 9 hours. \[Daniel ~collected~ \in ~1~ hour=\frac{ 40 }{ 9 }~bushels\] \[Both collected~\in~1~hour=\frac{ 40 }{ 4.95 }~bushels~of~apples.\] Paul collected in 1 hour\[=\frac{ 40 }{ 4.95 }-\frac{ 40 }{ 9 }=40\left( \frac{ 100 }{ 495 } -\frac{ 1 }{ 9 }\right)\] =?